# Muon beams towards muonium physics: progress and prospects

### v1

@@ IV.4 Precise measurement of antimatter gravity
The majority of the hydrogen atom’s rest mass is caused by the strong interaction within its nucleus. In contrast, hydrogen-like purely leptonic bound states, such as positronium () and muonium (), derive almost all of their mass from the leptons, making them more suitable candidates for gravity measurement. Moreover, muonium is currently the most accessible bound-state atom containing a second-generation lepton. Therefore, muonium has been considered as a promising subject for investigations of antimatter gravity. All current proposals for muonium gravity measurements are based on atom interferometry [Kirch:2014mna, Phillips:2018twx, MAGE:2018wxk, Soter:2021xuf]. The experimental principle can be summarized as follows (as also can be seen in Fig. 23): (1) muons are stopped in a target and emitted into vacuum as muonium; (2) muoniums pass through an interferometer consisting of three gratings with a period of nm; (3) under gravity, muoniums fall freely through a vertical displacement in flight, causing a vertical phase shift of the interference fringes; (4) at the end, coincident detection of the Michel positron and the low-energy orbital electron is performed to determine the muonium transmission rate. This experimental method demands a very high-quality muonium beam. As a result, current proposals tend to use SFHe, as mentioned in Section IV.1, instead of silica aerogel as the muonium formation target. There are two experiments aimed at the behavior of muonium under gravity, known as the Muonium Antimatter Gravity Experiment (MAGE) [MAGE:2018wxk] and LEptons in Muonium Interacting with Gravity (LEMING) [Soter:2021xuf], which are still in the technical design and demonstration stage.
@@ V Applications of muon beams
As we discussed in the previous section, muons are widely used in scientific research due to their unique physical properties. Meanwhile, muon-related applications have been developed with prosperity in recent years. Before the advent of accelerator muon beamlines, muon-related applications were mainly restricted to muongraphy using cosmic-ray muons. Nowadays, advanced muon beamlines provide an environment to advance muon-related application technologies and broaden the scope of muon science. Since the application technologies’ principles and setups are highly dependent on the charge of the muon beam, the following section provides an overview of the technologies for positive and negative beamlines, respectively.
@@ V.1.1 Basic concepts of SR
After establishing a muon beamline with promising polarization characteristics, the next step is to inject the muon beam into the sample under investigation. The interaction processes that occur during muon injection into the material are illustrated in Fig. 24. Assuming an initial muon momentum of , the muon is rapidly slowed down after entering the material due to multiple scattering with electrons. This scattering occurs on a timescale of approximately 1 ns and reduces the kinetic energy of muon to about 2–3 keV. At this point, the muon may form a bound state with electrons or atoms in the material. For positive muons (), it will capture an electron to form muonium (), while the remaining particles will remain unbound and free. The free and muonium typically retain kinetic energies in the -. In the case of negative muons (), some will interact with atomic electrons to form muonic atoms. In contrast, others will remain free and stable, with kinetic energies of a few keV. For free , depolarization effects are negligible. However, depolarization becomes significantly more pronounced when muons form bound states such as muonium or muonic atoms. For the muonic atom consisted of a muon and a non-zero nucleon, the muon will suffer from a strong depolarization effect, leading to the loss of most of the initial polarization. In contrast, the muon in muonium is only slightly depolarized due to the hyperfine interaction between the electron and the muon. Therefore, keeps a stronger polarization than when passing through the material, which serves a better role for the SR technique construction.
@@
where is the angle between the muon spin vector and the Michel electron momentum . The coefficient originates from integrating over the Michel electron energy spectrum. The angular distribution relative to and is plotted in Fig. 25.
Commonly, an array of detectors surrounding the sample measures the number of decay positrons arriving at different locations over time to reconstruct the polarization evolution. The typical configuration of a SR spectrometer is illustrated in Fig. 26. The detector modules are typically arranged symmetrically around the sample to capture positrons emitting in various directions. The detectors are usually scintillator counters coupled to photomultiplier tubes (PMTs) for high-efficiency positron detection. Recently, silicon photomultipliers (SiPMs) have also been employed as alternatives to PMTs due to their compact size and insensitivity to magnetic fields [Amato:2017gps, Kojima:2014qwv, Yang:2025lae]. By applying different external fields, the muon polarization reveals the magnetic properties of the samples. In the simple case of a zero magnetic field, the implanted muons’ spins do not precess. Consequently, the spatial distribution of the positrons remains static, reflecting only the initial spin polarization.
@@
where the is the muon lifetime, and the muon spin precession frequency is related to the local magnetic field via the muon gyromagnetic ratio :
@@
stands for the angle between the beamline direction and the direction from the sample pointing to the Michel electron detector. This transverse-field configuration is one of the most commonly used setups in SR experiments (TF-SR), as it directly connects the measurable precession frequency with the local magnetic field at the muon site. The typical TF-SR signal is shown in Fig. 27 (top left and bottom left). In other experimental options, such as longitudinal-field and zero-field SR, the relaxation of the polarization rather than its precession provides information about the static and dynamic characteristics of the local field distribution. Together, these complementary configurations allow SR to probe a wide range of magnetic phenomena in materials.
In muon spin relaxation measurements, no external magnetic field is applied (zero-field SR), or the external field is aligned parallel to the initial direction of the muon spin (longitudinal-field SR). The typical signals for these configurations are shown in Fig. 27 (top middle, top right, bottom middle, and bottom right). In these configurations, the muon spin does not precess coherently as in the transverse-field case. Instead, local magnetic fields at the muon sites cause the spin to relax gradually. These local fields arise from nearby atomic or nuclear moments in the material. When the local magnetic fields are randomly oriented, the longitudinal component of the muon spin exhibits time-dependent depolarization, described by
@@
Based on the relaxation measurements, the muon gradually depolarized under the local magnetic field. The muon spin resonance technique allowed active control of the spin state through an external radio-frequency (RF) magnetic field applied along the initial muon spin direction. The RF field can create Zeeman splitting, but it does not flip the muon spin directly. Instead, when the radio frequency matches the Zeeman energy gap, the muon spin will flip due to the hyperfine interaction. By combining precision, relaxation, and resonance measurements, researchers can fully probe the local magnetic properties of materials. The following sections describe applications of SR in condensed matter physics and materials science.
@@ V.1.2 SR technique in condensed matter physics
Beyond characterizing local contact fields, the application of SR under external magnetic fields is a powerful tool to investigate magnetic volume fractions and phase transitions. The magnetic field is typically a weak transverse field ranging from 5 to 10 . A distinct advantage of SR is its model-independent nature, which allows for the precise determination of the magnetic fraction without requiring prior assumptions about the magnetic structure of materials. In a characteristic temperature-dependent measurement, the experiment begins in the paramagnetic state, where the internal field at the muon site is negligible, and the spin precession is governed solely by the external field. As the material is cooled through a magnetic transition, the development of a spontaneous internal field leads to a reduction in the oscillation amplitude (or initial asymmetry) of the SR signal. By monitoring this amplitude across a range of temperatures, the evolution of the magnetic volume fraction can be quantitatively determined. Furthermore, the muon spin precession is robust against thermal fluctuations. Therefore, SR is suitable for probing magnetism at mK temperatures. Applications include the study of quantum phase transitions in Mott insulators [frandsen2016volume]. As shown in Fig. 28, SR spectroscopy measures the magnetic volume fraction at various temperatures. For instance, SR spectrometers equipped with pressure cells enable the study of magnetic transitions under high-pressure conditions [amato2004weak, bourdarot2005hidden]. Due to muons’ high sensitivity to weak magnetic fields, even organic magnets and systems with extremely small magnetic moments can be effectively investigated [le1993searching, blundell1995mu+, andreica2001magnetic].
With a characteristic time window of , SR is sensitive to magnetic fluctuations with correlation times spanning to . As shown in Fig. 29, this extensive range positions SR as an important complement to NMR and neutron scattering. Magnetic fluctuations are typically investigated through two approaches: varying the external field at constant temperature to map the fluctuation spectrum [aoki2003time, spehling2012magnetic], or varying the temperature at a fixed field to track dynamical evolution across phase transitions [kratzer2002musr]. In both cases, the measured muon depolarization rate provides a direct probe of the underlying fluctuation frequency. Beyond scaling behavior, SR can resolve fluctuation anisotropy by adjusting the sample orientation relative to the beamline [hartmann1991asymmetric]. More importantly, the technique can distinguish static from dynamic contributions to the local field, facilitating the study of systems with coexisting orders, such as heavy-fermion compounds [ran2019nearly, sundar2019coexistence] and uranium-based materials [de1995absence, keizer1999magnetism, amato2004weak].
@@
Consequently, when the external magnetic field is between the lower and upper critical fields (), magnetic flux enters the material in the form of quantized vortices. Each vortex carries a single quantum of magnetic flux and features a standard core. The SR technique effectively probes the internal magnetic field distribution within vortices, allowing for a detailed investigation of vortex lattice structures and dynamics in type II superconductors. Typically, the experimental measurement of vertices is performed using the transverse-field SR (TF-SR) technique, with a superconductor in the normal state under an external magnetic field. In this state, the local magnetic field at each muon site is essentially uniform, so all muon spins precess at the same frequency, producing a strong, coherent oscillation signal. As the sample cools into the superconducting state, vortices begin to form, and the magnetic field becomes spatially inhomogeneous due to the contribution from vortex cores. The muon spin precession frequency varies slightly depending on each muon’s local environment, leading to gradual depolarization of the signal. Upon further cooling well below the critical temperature, the magnetic inhomogeneity becomes more obvious, leading to rapid muon spin depolarization. The precession signal broadens significantly, reflecting the wide distribution of local magnetic fields associated with the vortex lattice. By analyzing the resulting muon precession spectrum, one can extract microscopic information such as the vortex density [Amato:1997zz].
@@ V.1.3 Muonium in SR
The applications of SR introduced above primarily rely on the properties of free muons. However, within matter, muons do not always remain in their free state, and form bound states instead. The bound muon state can either be muonium or a muonic atom. By detecting their decay products, one can investigate various properties at the atomic scale. In this section,e the fundamental mechanisms of muonium and its applications in SR will be introduced. According to the muonium basic properties which has been introduced in Section IV.3, the external magnetic field will lead to time-dependent muonium polarization in the material, which can be described as
where and is the hyperfine eigenstates of muonium system, is the transition frequency. Similar to the hyperfine splitting under the external magnetic field, the specific polarization time evolution depends on the strength and the direction of the external magnetic field. In a longitudinal magnetic field, the polarization is either along or opposite to the initial polarization direction. Then, the polarization time evolution can be expressed as
where is the dimensionless magnetic field, is the transition frequency between the muonium states and . The transition frequency can expressed as , where is the Larmor frequency. Therefore, if the external magnetic field is zero, the muonium polarization time evolution is only contributed by the Larmor precession . If the external magnetic field is perpendicular to the polarization direction, the muonium will be regularized. The regularization will lead to a non-oscillation part of the muonium polarization time evolution . If the external magnetic field is transverse, the muonium polarization time evolution is more complicated. In general, the polarization time evolution under the transverse field can be expressed as
@@
According to the Eq. 33, the polarization time evolution shows a beating spectrum with two frequencies and . If the external magnetic field is further increased, typically . In this case, the coefficients and . The polarization time evolution can be expressed as
In this case, the transition frequencies and . When , the sum of two observable frequency sums the hyperfine coupling of muonium by . If , the hyperfine coupling constant can be obtained by . Except for the hyperfine interaction, the muonium can also interact with the surrounding nuclei. The nuclear hyperfine interaction splits muonium energy levels, resulting in a complex frequency spectrum with many weak lines that cannot be resolved. If both the hyperfine interaction and the resulting muonium spin evolution can be treated as isotropic, the system is referred to as isolated muonium. This situation typically arises in nonconducting materials, where the captures an electron at an interstitial site. Such isolated muonium states are only found when the free-electron density around the muon is sufficiently low or when the muon occupies a highly symmetric lattice site. According to Cox’s work [cox2003shallow], the hyperfine constant inside solid state materials is smaller than that of free muonium. In elemental semiconductors, the hyperfine constant is typically about one-half of the free-muonium value. More generally, the hyperfine interaction need not be isotropic. In materials where the electronic environment is anisotropic, the hyperfine coupling is described not by a single constant but by a tensor . Such anitotropic muonium states have been reported in several semiconductors through the SR technique [cox2003shallow, cox2009muonium, patterson1988muonium].
Muonium can be a probe of hydrogen in physical chemistry. It behaves chemically like hydrogen, while being easier to generate and more stable under extreme conditions. Its light mass produces a obvious isotope effect, revealing reaction mechanisms such as tunneling-dominated addition reactions and inverse kinetic isotope effects in abstraction processes. This capability has been demonstrated in classic kinetic studies [Ghandi:2003muonium, Percival:2007Hkinetics], and more recently utilized to test quantum reaction rate theories and vibrational bonding phenomena that are inaccessible to heavier isotopes [Fleming2021PCCP]. Furthermore, SR enables the in situ formation and detection of radicals, providing access to their structures and dynamics through hyperfine interactions. This remains possible even at high temperatures, in supercritical fluids, or on very short timescales where conventional EPR techniques struggle. The method applies to a wide range of chemical environments, from gas-phase kinetics to soft matter.

### v2

@@ IV.4 Precise measurement of antimatter gravity
The majority of the hydrogen atom’s rest mass is caused by the strong interaction within its nucleus. In contrast, hydrogen-like purely leptonic bound states, such as positronium () and muonium (), derive almost all of their mass from the leptons, making them more suitable candidates for gravity measurement. Moreover, muonium is currently the most accessible bound-state atom containing a second-generation lepton. Therefore, muonium has been considered as a promising subject for investigations of antimatter gravity. All current proposals for muonium gravity measurements are based on atom interferometry [Kirch:2014mna, Phillips:2018twx, MAGE:2018wxk, Soter:2021xuf]. The experimental principle can be summarized as follows (as also can be seen in Fig.˜23): (1) muons are stopped in a target and emitted into vacuum as muonium; (2) muoniums pass through an interferometer consisting of three gratings with a period of nm; (3) under gravity, muoniums fall freely through a vertical displacement in flight, causing a vertical phase shift of the interference fringes; (4) at the end, coincident detection of the Michel positron and the low-energy orbital electron is performed to determine the muonium transmission rate. This experimental method demands a very high-quality muonium beam. As a result, current proposals tend to use SFHe, as mentioned in Section˜IV.1, instead of silica aerogel as the muonium formation target. There are two experiments aimed at the behavior of muonium under gravity, known as the Muonium Antimatter Gravity Experiment (MAGE) [MAGE:2018wxk] and LEptons in Muonium Interacting with Gravity (LEMING) [Soter:2021xuf], which are still in the technical design and demonstration stage.
@@ V Applications of muon beams
As we discussed in the previous section, muons are widely used in scientific research due to their unique physical properties. Meanwhile, muon-related applications have been developed with prosperity in recent years. Before the advent of accelerator muon beamlines, muon-related applications were mainly restricted to muography using cosmic-ray muons. Nowadays, advanced muon beamlines provide an environment to advance muon-related application technologies and broaden the scope of muon science. Since the application technologies’ principles and setups are highly dependent on the charge of the muon beam, the following section provides an overview of the technologies for positive and negative beamlines, respectively.
@@ V.1.1 Basic concepts of SR
After establishing a muon beamline with promising polarization characteristics, the next step is to inject the muon beam into the sample under investigation. The interaction processes that occur during muon injection into the material are illustrated in Fig.˜24. Assuming an initial muon momentum of , the muon is rapidly slowed down after entering the material due to multiple scattering with electrons. This scattering occurs on a timescale of approximately 1 ns and reduces the kinetic energy of muon to about 2–3 keV. At this point, the muon may form a bound state with electrons or atoms in the material. For positive muons (), they will capture an electron to form muonium (), while the remaining particles will remain unbound and free. The free and muonium typically retain kinetic energies of -. In the case of negative muons (), some will interact with atomic electrons to form muonic atoms. In contrast, others will remain free and stable, with kinetic energies of a few keV. For free , depolarization effects are negligible. However, depolarization becomes significantly more pronounced when muons form bound states such as muonium or muonic atoms. For the muonic atom consisting of a muon and a non-zero nucleon, the muon will undergo a strong depolarization effect, leading to the loss of most of the initial polarization. In contrast, the muon in muonium is only slightly depolarized due to the hyperfine interaction between the electron and the muon. Therefore, keeps a stronger polarization than when passing through the material, which plays a more significant role in the SR technique construction.
@@
where is the angle between the muon spin vector and the Michel electron momentum . The coefficient originates from integrating over the Michel electron energy spectrum. The angular distribution relative to and is plotted in Fig.˜25.
Commonly, an array of detectors surrounding the sample measures the number of decay positrons arriving at different locations over time to reconstruct the polarization evolution. The typical configuration of a SR spectrometer is illustrated in Fig.˜26. The detector modules are typically arranged symmetrically around the sample to capture positrons emitted in various directions. The detectors are usually scintillator counters coupled to photomultiplier tubes (PMTs) for high-efficiency positron detection. Recently, silicon photomultipliers (SiPMs) have also been employed as alternatives to PMTs due to their compact size and insensitivity to magnetic fields [Amato:2017gps, Kojima:2014qwv, Yang:2025lae]. By applying different external fields, the muon polarization reveals the magnetic properties of the samples. In the simple case of a zero magnetic field, the implanted muons’ spins do not precess. Consequently, the spatial distribution of the positrons remains static, reflecting only the initial spin polarization.
@@
where is the muon lifetime, and the muon spin precession frequency is related to the local magnetic field via the muon gyromagnetic ratio :
@@
stands for the angle between the beamline direction and the direction from the sample pointing to the Michel electron detector. This transverse-field configuration is one of the most commonly used setups in SR experiments (TF-SR), as it directly connects the measurable precession frequency with the local magnetic field at the muon site. The typical TF-SR signal is shown in Fig.˜27 (top left and bottom left). In other experimental options, such as longitudinal-field and zero-field SR, the relaxation of the polarization rather than its precession provides information about the static and dynamic characteristics of the local field distribution. Together, these complementary configurations allow SR to probe a wide range of magnetic phenomena in materials.
In muon spin relaxation measurements, no external magnetic field is applied (zero-field SR), or the external field is aligned parallel to the initial direction of the muon spin (longitudinal-field SR). The typical signals for these configurations are shown in Fig.˜27 (top middle, top right, bottom middle, and bottom right). In these configurations, the muon spin does not precess coherently as in the transverse-field case. Instead, local magnetic fields at the muon sites cause the spin to relax gradually. These local fields arise from nearby atomic or nuclear moments in the material. When the local magnetic fields are randomly oriented, the longitudinal component of the muon spin exhibits time-dependent depolarization, described by
@@
Based on the relaxation measurements, the muon gradually depolarizes under the local magnetic field. The muon spin resonance technique allowed active control of the spin state through an external radio-frequency (RF) magnetic field applied along the initial muon spin direction. The RF field can create Zeeman splitting, but it does not flip the muon spin directly. Instead, when the radio frequency matches the Zeeman energy gap, the muon spin will flip due to the hyperfine interaction. By combining precession, relaxation, and resonance measurements, researchers can fully probe the local magnetic properties of materials. The following sections describe applications of SR in condensed matter physics and materials science.
@@ V.1.2 SR technique in condensed matter physics
Beyond characterizing local contact fields, the application of SR under external magnetic fields is a powerful tool to investigate magnetic volume fractions and phase transitions. The magnetic field is typically a weak transverse field ranging from 5 to 10 . A distinct advantage of SR is its model-independent nature, which allows for the precise determination of the magnetic fraction without requiring prior assumptions about the magnetic structure of materials. In a characteristic temperature-dependent measurement, the experiment begins in the paramagnetic state, where the internal field at the muon site is negligible, and the spin precession is governed solely by the external field. As the material is cooled through a magnetic transition, the development of a spontaneous internal field leads to a reduction in the oscillation amplitude (or initial asymmetry) of the SR signal. By monitoring this amplitude across a range of temperatures, the evolution of the magnetic volume fraction can be quantitatively determined. Furthermore, the muon spin precession is robust against thermal fluctuations. Therefore, SR is suitable for probing magnetism at mK temperatures. Applications include the study of quantum phase transitions in Mott insulators [frandsen2016volume]. As shown in Fig.˜28, SR spectroscopy measures the magnetic volume fraction at various temperatures. For instance, SR spectrometers equipped with pressure cells enable the study of magnetic transitions under high-pressure conditions [amato2004weak, bourdarot2005hidden]. Due to muons’ high sensitivity to weak magnetic fields, even organic magnets and systems with extremely small magnetic moments can be effectively investigated [le1993searching, blundell1995mu+, andreica2001magnetic].
With a characteristic time window of , SR is sensitive to magnetic fluctuations with correlation times spanning to . As shown in Fig.˜29, this extensive range positions SR as an important complement to NMR and neutron scattering. Magnetic fluctuations are typically investigated through two approaches: varying the external field at constant temperature to map the fluctuation spectrum [aoki2003time, spehling2012magnetic], or varying the temperature at a fixed field to track dynamical evolution across phase transitions [kratzer2002musr]. In both cases, the measured muon depolarization rate provides a direct probe of the underlying fluctuation frequency. Beyond scaling behavior, SR can resolve fluctuation anisotropy by adjusting the sample orientation relative to the beamline [hartmann1991asymmetric]. More importantly, the technique can distinguish static from dynamic contributions to the local field, facilitating the study of systems with coexisting orders, such as heavy-fermion compounds [ran2019nearly, sundar2019coexistence] and uranium-based materials [de1995absence, keizer1999magnetism, amato2004weak].
@@
Consequently, when the external magnetic field is between the lower and upper critical fields (), magnetic flux enters the material in the form of quantized vortices. Each vortex carries a single quantum of magnetic flux and features a standard core. The SR technique effectively probes the internal magnetic field distribution within vortices, allowing for a detailed investigation of vortex lattice structures and dynamics in type II superconductors. Typically, the experimental measurement of vortices is performed using the transverse-field SR (TF-SR) technique, with a superconductor in the normal state under an external magnetic field. In this state, the local magnetic field at each muon site is essentially uniform, so all muon spins precess at the same frequency, producing a strong, coherent oscillation signal. As the sample cools into the superconducting state, vortices begin to form, and the magnetic field becomes spatially inhomogeneous due to the contribution from vortex cores. The muon spin precession frequency varies slightly depending on each muon’s local environment, leading to gradual depolarization of the signal. Upon further cooling well below the critical temperature, the magnetic inhomogeneity becomes more obvious, leading to rapid muon spin depolarization. The precession signal broadens significantly, reflecting the wide distribution of local magnetic fields associated with the vortex lattice. By analyzing the resulting muon precession spectrum, one can extract microscopic information such as the vortex density [Amato:1997zz].
@@ V.1.3 Muonium in SR
The applications of SR introduced above primarily rely on the properties of free muons. However, within matter, muons do not always remain in their free state, and form bound states instead. The bound muon state can either be muonium or a muonic atom. By detecting their decay products, one can investigate various properties at the atomic scale. In this section, the fundamental mechanisms of muonium and its applications in SR will be introduced. According to the muonium basic properties which have been introduced in Section˜IV.3, the external magnetic field will lead to time-dependent muonium polarization in the material, which can be described as
where and are the hyperfine eigenstates of muonium system, is the transition frequency. Similar to the hyperfine splitting under the external magnetic field, the specific polarization time evolution depends on the strength and the direction of the external magnetic field. In a longitudinal magnetic field, the polarization is either along or opposite to the initial polarization direction. Then, the polarization time evolution can be expressed as
where is the dimensionless magnetic field, is the transition frequency between the muonium states and . The transition frequency can be expressed as , where is the Larmor frequency. Therefore, if the external magnetic field is zero, the muonium polarization time evolution is only contributed by the Larmor precession . If the external magnetic field is perpendicular to the polarization direction, the muonium will be regularized. The regularization will lead to a non-oscillation part of the muonium polarization time evolution . If the external magnetic field is transverse, the muonium polarization time evolution is more complicated. In general, the polarization time evolution under the transverse field can be expressed as
@@
According to the Eq.˜33, the polarization time evolution shows a beating spectrum with two frequencies and . If the external magnetic field is further increased, typically . In this case, the coefficients and . The polarization time evolution can be expressed as
In this case, the transition frequencies and . When , the sum of two observable frequencies yields the hyperfine coupling of muonium by . If , the hyperfine coupling constant can be obtained by . Except for the hyperfine interaction, the muonium can also interact with the surrounding nuclei. The nuclear hyperfine interaction splits muonium energy levels, resulting in a complex frequency spectrum with many weak lines that cannot be resolved. If both the hyperfine interaction and the resulting muonium spin evolution can be treated as isotropic, the system is referred to as isolated muonium. This situation typically arises in nonconducting materials, where the captures an electron at an interstitial site. Such isolated muonium states are only found when the free-electron density around the muon is sufficiently low or when the muon occupies a highly symmetric lattice site. According to Cox’s work [cox2003shallow], the hyperfine constant inside solid state materials is smaller than that of free muonium. In elemental semiconductors, the hyperfine constant is typically about one-half of the free-muonium value. More generally, the hyperfine interaction need not be isotropic. In materials where the electronic environment is anisotropic, the hyperfine coupling is described not by a single constant but by a tensor . Such anisotropic muonium states have been reported in several semiconductors through the SR technique [cox2003shallow, cox2009muonium, patterson1988muonium].
Muonium can be a probe of hydrogen in physical chemistry. It behaves chemically like hydrogen, while being easier to generate and more stable under extreme conditions. Its light mass produces an obvious isotope effect, revealing reaction mechanisms such as tunneling-dominated addition reactions and inverse kinetic isotope effects in abstraction processes. This capability has been demonstrated in classic kinetic studies [Ghandi:2003muonium, Percival:2007Hkinetics], and more recently utilized to test quantum reaction rate theories and vibrational bonding phenomena that are inaccessible to heavier isotopes [Fleming2021PCCP]. Furthermore, SR enables the in situ formation and detection of radicals, providing access to their structures and dynamics through hyperfine interactions. This remains possible even at high temperatures, in supercritical fluids, or on very short timescales where conventional EPR techniques struggle. The method applies to a wide range of chemical environments, from gas-phase kinetics to soft matter.